### v2

# NVILA: Efficient Frontier Visual Language Models

Abstract
Abstract:   Visual language models (VLMs) have made significant advances in accuracy in recent years. However, their efficiency has received much less attention. This paper introduces NVILA, a family of open VLMs designed to optimize both efficiency and accuracy. Building on top of VILA, we improve its model architecture by first scaling up the spatial and temporal resolutions, and then compressing visual tokens. This “scale-then-compress” approach enables NVILA to efficiently process high-resolution images and long videos. We also conduct a systematic investigation to enhance the efficiency of NVILA throughout its entire lifecycle, from training to deployment. NVILA matches or surpasses the accuracy of many leading open and proprietary VLMs across a wide range of image and video benchmarks. At the same time, it reduces training costs by 1.9-5.1$\times$, prefilling latency by 1.6-2.2$\times$, and decoding latency by 1.2-2.8$\times$. We make our code and models available to facilitate reproducibility.
Links:Code (on GitHub) | Models (on Hugging Face) | Demo | Subscribe
Figure 1: NVILA – Efficient Frontier VLMs. (a) NVILA trains image and video models 5.1$\times$ and 1.9$\times$ faster, respectively, than LLaVA-OneVision (OV), which is the only baseline model with publicly disclosed training costs. (b) Against Qwen2-VL, NVILA achieves a 1.6–2.2$\times$ measured speedup in the pre-filling stage and a 1.2–2.8$\times$ speedup during the decoding stage. (c) NVILA’s efficiency is achieved without compromising accuracy; in fact, it delivers comparable or even superior accuracy across image and video benchmarks. All models in this table have 8B parameters. Training time in (a) is measured using NVIDIA H100 GPUs, while inference speed in (b) is measured using a single NVIDIA GeForce RTX 4090 GPU. Accuracy numbers in (c) are normalized relative to the highest score for each benchmark.
Figure 2: Qualitative examples.
Figure 3: Model architecture.
1 Introduction
Visual language models (VLMs) have shown remarkable abilities in processing and integrating both visual and textual information, enabling advanced vision-language interactions and dialogues. In recent years, the research community has made tremendous progress in enhancing their accuracy [1, 2, 3, 4, 5] and broadening their applications across diverse domains, including robotics [6, 7, 8], autonomous driving [9], and medical applications [10, 11]. However, there has been much less focus on improving their efficiency.
VLMs are expensive across multiple dimensions. First, training a VLM is time-consuming. For example, training a state-of-the-art 7B VLM [4] can take up to 400 GPU days, let alone even larger models. This creates a significant entry barrier for researchers. Second, VLMs often require adaptation when applied to specialized domains (e.g., medical imaging), but fine-tuning a VLM is memory-intensive. For example, fully fine-tuning a 7B VLM can require over 64GB of GPU memory, far beyond the available memory of most consumer-level GPUs. Finally, VLMs are often deployed in edge applications with limited computational budget (e.g., laptops, robots), so deploying a VLM is resource-constrained. Addressing these challenges requires a systematic solution to improve VLM efficiency across all these dimensions.
In this paper, we introduce NVILA, a family of open VLMs designed to optimize both efficiency and accuracy. Building on VILA [2], we improve its model architecture by first scaling up the spatial and temporal resolution, followed by compressing visual tokens. “Scaling” preserves more details from visual inputs, raising the accuracy upper bound, while “compression” squeezes visual information to fewer tokens, improving computational efficiency. This “scale-then-compress” strategy allows NVILA to process high-resolution images and long videos both effectively and efficiently. In addition, we conduct a systematic study to optimize the efficiency of NVILA throughout its entire lifecycle, including training, fine-tuning, and deployment.
Thanks to these innovations, NVILA is efficient and accurate. It reduces training costs by 1.9-5.1$\times$, prefilling latency by 1.6–2.2$\times$, and decoding latency by 1.2–2.8$\times$. It also matches or surpasses the accuracy of leading open VLMs [5, 3, 2] and proprietary VLMs [12, 13] across a wide range of image and video benchmarks. Furthermore, NVILA enables new capabilities including temporal localization, robotic navigation, and medical imaging. We release our code and models to support full reproducibility. We hope our work will inspire further research on efficient VLMs.
2 Approach
In this section, we begin by designing an efficient model architecture for NVILA, first by scaling up spatial and temporal resolutions, and then by compressing the visual tokens. Next, we present strategies to improve NVILA’s efficiency across its entire lifecycle—from training and fine-tuning to deployment. Unless otherwise specified, all analysis in this section will be based on the 8B model.
2.1 Efficient Model Architecture
We build NVILA on top of VILA [2]. As in Figure 3, it is an auto-regressive VLM composed of three components: a visual encoder that extracts features from visual inputs (e.g., images, videos); a projector that aligns embeddings across visual and language modalities; and a token processor, typically instantiated with a LLM, which takes both visual and language tokens as input and outputs language tokens. Specifically, NVILA uses SigLIP [14] as its vision encoder, a two-layer MLP as its projector, and Qwen2 [15] of different sizes as its token processor.
The original VILA has very limited spatial and temporal resolutions: i.e., it resizes all images to 448$\times$448, regardless of their original size or aspect ratio, and samples up to 14 frames from videos***This is the configuration for VILA-1.5 40B. Their other variants, such as VILA-1.5 3B, only use 384$\times$384 resolution and 8 frames.. Both spatial resizing and temporal sampling will introduce significant loss of information, limiting the model’s capability to effectively process larger images and longer videos. This can also be observed in Table 8 and Table 9, where VILA lags behind leading VLMs, especially on text-heavy and long-video benchmarks.
In this paper, we advocate for the “scale-then-compress” paradigm, where we first scale up the spatial/temporal resolutions to improve accuracy, and we then compress the visual tokens to improve efficiency. Scaling resolutions up improves the performance ceiling, but doing so alone will significantly increase the computational cost. For example, doubling the resolution will double the number of visual tokens, which will increase both training and inference costs by more than 2$\times$, as self-attention scales quadratically with the number of tokens. We can then cut this cost down by compressing spatial/temporal tokens. Compressed visual tokens have a higher information density, allowing us to preserve or even improve spatial and temporal details with fewer total tokens.
Table 1: Spatial “scale-then-compress”. Increasing the spatial resolution with Dynamic-S2 can greatly improve the model’s accuracy, particularly on text-heavy benchmarks. Compressing the visual tokens with spatial pooling can effectively reduce both the number of tiles and tokens per tile, with moderate accuracy loss. This loss can be further reduced by adding an additional visual encoder pre-training (VEP) stage. In this and following tables, “IM-10” refers to the average validation scores from the 10 benchmarks listed in Table 8.
Spatial Pooling #Tokens/Tile #Tiles/Image AI2D DocVQA TextVQA IM-10
Baseline (VILA-1.5) 2$\times$2 256 (=16$\times$16) 1 87.0 61.3 67.5 61.2
Scale (Dynamic-S2) 2$\times$2 256 (=16$\times$16) 9-12 90.1 91.1 77.0 71.5
Scale + Compress 3$\times$3 121 (=11$\times$11) 1-12 87.4 82.3 74.1 67.1
Scale + Compress + VEP 3$\times$3 121 (=11$\times$11) 1-12 89.8 88.8 76.1 70.8
Alternative Designs
TokenLearner 121 1-12 90.0 86.5 75.6 69.8
Perceiver Resampler 121 1-12 76.8 71.8 65.3 59.4
Table 2: Temporal “scale-then-compress”. Scaling up the temporal resolution can improve the model’s video understanding performance. Compressing the visual tokens with temporal averaging can effectively reduce the number of tokens with only a marginal accuracy drop.
#Frames Temporal Pooling #Tokens/Video Video-MME (w/o sub.)
Short Medium Long Overall
Baseline (VILA-1.5) 8 1$\times$ 2048 (=162$\times$8) 65.4 53.8 47.7 55.7
Scale 32 1$\times$ 8192 (=162$\times$32) 73.2 58.9 50.9 61.0
Scale + Compress 32 4$\times$ 2048 (=162$\times$32/4) 73.7 56.7 50.0 60.1
Scale + Compress 256 8$\times$ 8192 (=162$\times$256/8) 75.0 62.2 54.8 64.0
2.1.1 Spatial “Scale-Then-Compress”
For spatial scaling, it is very natural to directly increase the image resolution of the vision encoder, for example, to 896$\times$896. While this may improve performance, applying a uniformly high resolution to all images would be inefficient, especially for smaller images that do not require extensive detail. To address this, we apply S2 [16] to efficiently extract multi-scale high-resolution features with image tiling. For example, given a vision encoder pre-trained at 4482 resolution and an input image with any size, S2 first resizes the image into multiple scales (e.g., 4482, 8962, 13442), and for each scale, it splits the image into tiles of 4482. Each tile is then individually processed by the encoder. The feature maps of each tile from the same scale are stitched back together into the feature map of the whole image at that scale. Finally, feature maps from different scales are interpolated into the same size and concatenated on the channel dimension.
S2 always resizes images into square, regardless of the original aspect ratio. This can cause distortion, particularly for images that are either tall and narrow or short and wide. To address this, we propose Dynamic-S2, which adaptively processes images with varying aspect ratios. Dynamic-S2 follows the approach of S2 but, at the largest image scale, instead of resizing to a square, it adjusts the image dimensions to the closest size that maintains the original aspect ratio and is divisible by 4482 tiles. This is inspired by the dynamic resolution strategy in InternVL [17]. After processing the tiles, the feature maps from all scales are interpolated to match the size of the largest scale and concatenated.
Equipped with Dynamic-S2, the model benefits from high-resolution information from the image, resulting in a up to 30% accuracy improvements on text-heavy benchmarks (Table 1). Our goal, then, shifts to compressing the spatial tokens. VILA [2] finds that applying a simple 2$\times$2 spatial-to-channel (STC) reshape can reduce the token count by a factor of 4 without sacrificing accuracy. However, pushing this further results in a notable drop in performance: i.e., a nearly 10% decrease in accuracy on DocQA, when reducing the number of minimal tiles and increasing the STC to 3$\times$3. We hypothesize that more aggressive reductions make the projector significantly harder to train. To address this, we introduce an additional visual encoder pre-training stage to jointly tune the vision encoder and projectors. This helps recover most of the accuracy loss from spatial token reduction, achieving a 2.4$\times$ speedup in both training and inference.
There are many alternative designs for spatial token compression, such as TokenLearner from RT-1 [6] and Perceiver Resampler from MiniCPM-V [18]. However, with the same token reduction ratio, these learnable compression methods surprisingly do not perform better than the simple spatial-to-channel design, even with an additional stage 1.5. We believe this is more of an optimization problem and is beyond the scope of this paper.
2.1.2 Temporal “Scale-Then-Compress”
For temporal scaling, we simply increase the number of uniformly sampled frames from the input video. Following previous methods [19], we train the model with additional video-supervised fine-tuning (SFT) to extend its capability to process more frames. From Table 9, extending the number of frames from 8 to 32 can increase the model’s accuracy on Video-MME by more than 5%. However, this will also increase the number of visual tokens by 4$\times$.
Similar to spatial token compression, we will then reduce these visual tokens. Since there is intrinsic temporal continuity in the video, we adopt temporal averaging [20] for compression, which first partitions the frames into groups and then temporally pools visual tokens within each group. This will reduce temporal redundancy (since consecutive frames often contain similar information) while still retaining important spatiotemporal information. Empirically, compressing the visual tokens by 4$\times$ leads to an acceptable accuracy drop. When compared to the original baseline with the same number of tokens, the first scaled and then expanded result costs almost the sameWe will need to run visual encoder for more frames, but this is usually not the runtime bottleneck., but has much higher accuracy. We have also used this approach to further scale the number of frames and the compression ratio, leading to a state-of-the-art 7B model on this benchmark (see Table 9).
2.2 Efficient Training
While state-of-the-art VLMs boast impressive capabilities, training such a VLM is often costly and compute-intensive. This section explores system-algorithm co-design to enable efficient VLM training. On the algorithm front, we examine a novel unsupervised dataset pruning method to streamline training data. At the system level, we investigate FP8 mixed precision for acceleration.
Figure 4: Dataset pruning. DeltaLoss visualizations in NVILA training: Left, Middle, and Right sections show examples that are too easy, distracting, and helpful for training, respectively.
2.2.1 Dataset Pruning
In order to improve model accuracy, previous work [21, 4, 22] kept grabbing high quality SFT datasets from various sources and can show improvement on Benchmark scores. However, not all data contributes equally to the model and continuous growth of datasets lead to much redundancy. In NVILA, we follow the “Scale-Then-Compress” concept to first increase our SFT dataset mixture and then trying to compress the dataset. However, selecting high-quality examples from various sources is challenging. While there have been explorations of vision inputs [23, 24, 25] and text-only inputs [26, 27, 28], few studies have addressed this problem in VLM training, where images and texts are mixed during training. NVILA’s training involves more than 100M data, making it necessary to prune the training set while maintaining accuracy.
Inspired by recent works in knowledge distillation [29], we leverage DeltaLoss to score the training set:
$D^{\prime}=\bigcup_{i=1}^{K}\text{top-}K\left\{\left.\log\frac{p_{\text{large}}(x)}{p_{\text{small}}(x)}\right|x\in D_{i}\right\},$ (1)
where $D_{i}$ is the $i$-th subset of the full fine-tuning datasets and $D^{\prime}$ is the pruned training set. $p_{\text{large}}(x)$ and $p_{\text{small}}(x)$ are the output probabilities on the answer tokens. The main motivation is to filter out examples that are either too easy or too hard. To elaborate,
  • If both answer correctly or wrongly, $\log\frac{p_{\text{large}}(x)}{p_{\text{small}}(x)}$ is close to 0.
  • When the small model answers correctly but the large model fails, $\log\frac{p_{\text{large}}(x)}{p_{\text{small}}(x)}$ becomes negative, suggesting these examples tend to distract learning and will eventually be forgotten by a more powerful model.
  • When the small model answers incorrectly but the large model solves it, $\log\frac{p_{\text{large}}(x)}{p_{\text{small}}(x)}$ is positive, suggesting these examples provide strong supervision, as challenging for small models but learnable by larger ones.
Thereby we can apply DeltaLoss to each sub-dataset and prune the training set with different ratios.
Table 3: Dataset pruning on NVILA Recipe. DeltaLoss consistently rivals other data selection methods and shows negligible performance drop when pruning 50% of data.
Method IM-10 MMMU DocVQA TextVQA
100% (baseline) 75.6 48.0 90.1 78.8
50%
DeltaLoss [29] 75.5 48.1 89.7 78.4
Cluster Pruning 74.5 47.8 88.3 77.0
Random Pruning 74.0 47.6 87.1 76.6
30%
DeltaLoss [29] 74.0 47.8 87.9 76.4
Cluster Pruning 73.5 47.7 84.1 76.0
Random Pruning 73.1 47.7 82.9 75.6
10%
DeltaLoss [29] 72.4 47.1 84.4 74.5
Cluster Pruning 72.2 47.4 79.6 73.2
Random Pruning 72.0 47.0 77.3 72.6
To evaluate the data pruning criterion, we compare DeltaLoss and the random pruning baseline in Table 3. For random pruning, data is randomly selected and we run the results three times and report the average. For cluster pruning, we apply k-means clustering with siglip features and prune the data evenly across each centroid. Our experiments report the average performance across 10 benchmarks, with a focus on key tasks to demonstrate the method’s effectiveness. We examine three pruning threshold 10%, 30% and 50% and notice that DeltaLoss consistently outperforms the random baseline, especially on the GQA and DocVQA tasks the random pruning shows a significant performance degradation while DeltaLoss stays accurate. We notice 50% is a relatively safe threshold where the average score maintains competitive while the training can be speedup by 2$\times$. Thus we set the threshold to 50% for later experiments.
We examine the impact of data pruning on newly added datasets. We incorporate varying percentages of pixmo data [30] into the NVILA training set. In Table LABEL:tab:dataset_prune_pixmo, we observe that directly combining pixmo data with the NVILA training set decreases performance on DocVQA and TextVQA benchmarks, while only improving MMMU scores. This suggests that aggressively increasing the training set size may actually hurt performance. By applying deltaloss to prune the training data—filtering out examples that are either too easy or too hard—we find that models trained with the pruned molmo dataset show general improvements in experimental results.
2.2.2 FP8 Training
Table 4: FP8 training. FP8 accelerates the training of NVILA while maintaining the accuracy, especially when gradient checkpointing (GC) is not enabled. In this table, the throughput results are obtained with the maximum achievable batch size (BS) on 64 H100 GPUs. Video-MME results come from an 8-frame setting and with subtitle information.
GC BS Throughput MMMU Video-MME
BF16 4 199.2 (1.0$\times$) 47.9 52.9
FP8 16 390.1 (2.0$\times$) 47.0 53.0
BF16 30 491.7 (2.5$\times$) 47.8 53.1
FP8 36 579.9 (2.9$\times$) 47.7 53.0
FP16 [31] and BF16 [32] are standard precisions for model training, since they offer acceleration without accuracy loss, supported natively by NVIDIA GPUs. With the advent of the NVIDIA Hopper and Blackwell architectures, new GPUs (such as H100 and B200) now provide native support for FP8, which has emerged as a promising precision due to its potential for larger computational and memory efficiency.
Many researchers have already applied FP8 to LLM training. NVIDIA’s Transformer Engine performs matrix multiplications (GEMM) in FP8 precision, resulting in faster training speeds. FP8-LM [33] builds upon this by also quantizing the gradients, weight master copy, and first-order momentum into FP8, thereby reducing communication overhead and memory footprint. COAT [34] further compresses activations and the optimizer’s second-order momentum to enhance memory efficiency while maintaining accuracy.
In this paper, we borrow the FP8 implementation from COAT [34] to accelerate the training of NVILA. One key difference between LLM and VLM training workloads lies in the variability of sequence lengths across batches. In LLM training, samples generally have uniform lengths, and increasing the batch size beyond a certain point has minimal effect on training throughput. However, in VLM training, samples can vary significantly in length: video samples may require tens of thousands of tokens, image samples may need hundreds, and text-only samples require far fewer. As a result, workloads with fewer tokens are generally underutilized and can benefit greatly from increasing the batch size. As shown in Table 4, applying FP8 to both weights and activations allows NVILA to increase the batch size from 4 to 16, resulting in a 2$\times$ speedup. When gradient checkpointing is enabled, quantizing activations becomes less essential. Instead, we integrate the cross-entropy kernel from Liger [35] to reduce peak memory usage due to Qwen’s large vocabulary size. In this case, FP8 training can still provide a 1.2$\times$ speedup compared to BF16 training.
Table 5: Fine-tuning recipe. Our recommendation is to tune the LLM with either LoRA or QLoRA and to tune ViT’s layer normalization (LN) layers with a much smaller learning rate. This setup achieves competitive accuracy and is also the most memory- and compute-efficient. All experiments use a batch size of 1 with gradient checkpointing disabled, and throughput is measured on a single NVIDIA A100 80GB GPU. For settings with {1,5,10,50}, we select the learning rate ratio from this set that gives the best results for each benchmark. “FT-5” refers to the average accuracy across AITZ [36], ALFRED [37], nuScenes [38], PathVQA [39], and Widget Caption [40].
ViT LLM Memory (GB) Throughput (iter/s) $\texttt{LR}_{\texttt{LLM}}/\texttt{LR}_{\texttt{ViT}}$ Accuracy (FT-5)
LoRA LoRA 20.1 3.4 1 69.2
{1,5,10,50} 71.8
LN LoRA 19.2 4.5 1 63.5
{1,5,10,50} 71.4
FT LoRA 21.9 4.2 1 64.0
{1,5,10,50} 70.1
LoRA QLoRA 11.1 2.6 1 63.0
{1,5,10,50} 70.8
LN QLoRA 10.2 3.1 1 62.7
{1,5,10,50} 70.9
FT FT 63.5 6.1 1 77.7
2.3 Efficient Fine-Tuning
Once a foundation VLM is trained, domain-specific fine-tuning is needed to adapt the model for specialized tasks or domains. While fine-tuning effectively improves domain-specific vocabulary and concepts, conventional Parameter Efficient Fine-Tuning has been focusing on LLM and text-related tasks, but how to best fine-tune a VLM remains less explored. In NVILA, we find that (i) The learning rate should be set differently for ViT and LLMs (ii) The tuning parts should be chosen dependently for different downstream tasks.
When fine-tuning the vision encoder (ViT) and language model (LLM) together using PEFT methods, we observe that the learning rate should be set differently for VE and LLM: the learning rate for the ViT part will be 5-50$\times$ smaller than that for the LLM part. On the other hand, we also observe that fine-tuning the vision encoder with Layernorm can achieve comparable performance as LoRA (Table. 5) while being more computationally efficient: it can reduce the training time by 25% compared to applying LoRA for the vision encoder. With the curated configuration setup, NVILA can be quickly fine-tuned to various downstream tasks under 24 GB memory with on-par performance.
Table 6: Quantization recipe. While W4A16 quantization on LLM backbone may introduce small accuracy drop, W8A8 quantization on ViT is nearly lossless.
ViT LLM AI2D MMMU VideoMME TTFT (s)
FP16 FP16 91.0 50.7 63.9 0.90
FP16 W4A16 90.9 49.2 62.0 0.77
W8A8 W4A16 90.9 49.3 62.1 0.65
2.4 Efficient Deployment
VLMs are often integrated in edge applications as robotic where computational budget is tight. In this section, we will introduce our inference engine with quantization to accelerate the deployment.
We develop a specialized inference engine with quantization techniques to efficiently deploy NVILA. The inference process is divided into two phases: prefilling and decoding. In the compute-bounded prefilling stage, we first apply token compression techniques (Section 2.1) to reduce the inference workload for LLM backbone, after which the vision tower becomes the primary bottleneck, accounting for over 90% of the prefilling latency. To tackle this, we implement W8A8 quantization for the vision tower to reduce NVILA’s Time-To-First-Token (TTFT) in this compute-bounded stage. For the memory-bounded decoding stage, we follow AWQ [41] for W4A16 quantization of the LLM backbone to accelerate. We further optimize the original AWQ implementation by introducing FP16 accumulation to the W4A16 GEMM kernels, resulting to a total 1.7$\times$ kernel speedup without compromising accuracy. We attach a detailed comparison in Figure. 5.
3 Experiments
3.1 Training Details
Table 7: Training recipe. Building upon VILA, we introduce two additional stages for NVILA: Stage 2, which focuses on pre-training the visual encoder to reduce performance loss due to spatial token compression, and Stage 5, which focuses on video instruction tuning to improve the model’s long video capability.
Visual Encoder Projector Token Processor
(ViT) (MLP) (LLM) LR
Initial from [14] random from [15]
Stage 1 frozen trainable frozen 1$\times$10-3
Stage 2 trainable trainable frozen 5$\times$10-5
Stage 3 frozen trainable trainable 5$\times$10-5
Stage 4 trainable trainable trainable 2$\times$10-5
Stage 5 trainable trainable trainable 2$\times$10-5
We follow a five-stage pipeline to train NVILA: (1) projector initialization, (2) visual encoder pre-training, (3) token processor pre-training, (4) image instruction-tuning, and (5) video instruction-tuning. Among them, Stages 1, 3, and 4 are also included in VILA training. The additional Stage 2 is used to recover the accuracy loss due to spatial token compression (as in Table 1), and the additional Stage 5 is helpful for extending the model’s long video understanding capability. We provide the detailed training recipe in Table 7 and data recipe in Table A1.
Our implementation is built upon PyTorch 2.3.0 [42, 43] and Transformers 4.46.0 [44]. We use DeepSpeed 0.9.5 [45] to shard large models across devices and use gradient checkpointing to reduce memory usage. We adopt FlashAttention-2 [46] to accelerate training in both the LLM and visual encoder. We also implement functional-preserving, on-the-fly sequence packing to fuse samples with different lengths, which leads to an around 30% speedup. We train all models using 128 NVIDIA H100 GPUs with a global batch size of 2048 across all stages. All optimizations are carried out using AdamW with no weight decay. We adopt a cosine learning rate decay schedule with a linear warm-up for the first 3% of the schedule. The initial learning rate varies across stages, as detailed in Table 7.
3.2 Accuracy Results
Table 8: Image benchmarks. We mark the best performance bold and the second-best underlined.
AI2D ChartQA DocVQA InfoVQA MathVista MMMU Real- WorldQA SEED TextVQA VQAv2
test test test test testmini val test pro image val testdev
GPT-4o 94.2 85.7 92.8 79.2 63.8 69.1 64.7 51.9 75.4 76.2 77.4 78.7
Claude 3.5 Sonnet 94.7 90.8 85.2 74.3 67.7 68.3 63.7 51.5 60.1 74.1 70.7
Gemini 1.5 Pro 94.4 87.2 93.1 81.0 63.9 62.2 57.6 43.5 70.4 78.7 80.2
LLaVA-1.5 7B 55.5 17.8 28.1 25.8 25.6 35.7 54.8 66.1 58.2 78.5
VILA-1.5 8B 76.6 52.7 40.6 25.9 36.7 38.6 32.7 52.7 73.8 68.5 83.0
Cambrian-1 8B 73.0 73.3 77.8 41.6 49.0 42.7 64.2 74.7 71.7 81.2
Florence-VL 8B 74.2 74.7 84.9 51.7 55.5 43.7 64.2 74.9 74.2 84.7
LLaVA-OneVision 8B 81.4 80.0 87.5 68.8 63.2 48.8 42.8 24.1 66.3 75.4 78.3 84.0
Llama 3.2 11B 91.9 83.4 88.4 51.5 50.7 75.2
InternVL2 8B 83.8 83.3 91.6 74.8 58.3 51.2 42.6 29.0 64.2 76.2 77.4 76.7
Qwen2-VL 8B 83.0 83.0 94.5 76.5 58.2 54.1 46.6 30.5 70.1 76.0 84.3 82.9
NVILA-Lite 8B 91.0 84.8 91.7 67.9 64.5 50.7 45.7 26.5 65.6 76.3 78.1 85.0
NVILA 8B 92.3 86.1 93.7 70.7 65.4 49.9 44.4 27.8 68.6 76.5 80.1 85.4
LLaVA-1.5 13B 61.1 18.2 30.3 29.4 27.7 37.0 55.3 68.2 61.3 80.0
VILA-1.5 13B 79.9 59.5 58.6 30.4 42.7 37.9 33.6 57.5 72.6 65.0 82.8
Cambrian-1 13B 73.6 73.8 76.8 48.0 40.0 63.0 74.4 72.8
Pixtral 12B 79.0 81.8 90.7 50.8 58.0 52.5 65.4 75.7 80.2
NVILA-Lite 15B 92.0 81.8 90.6 69.3 61.7 58.7 51.8 33.7 67.1 75.6 77.3 83.7
NVILA 15B 94.1 86.9 94.0 73.5 66.1 56.7 51.8 33.8 69.5 76.6 80.0 84.8
LLaVA-NeXT 34B 46.5 48.1 44.5 22.9 75.9 69.5 83.7
Cambrian-1 34B 79.7 75.6 75.5 46.0 53.2 49.7 67.8 75.3 76.7 83.8
VILA-1.5 40B 88.9 67.8 58.6 38.4 49.3 51.9 46.9 25.0 60.8 69.1 73.6 84.3
InternVL2 40B 87.1 86.2 93.9 78.7 63.7 55.2 47.4 34.2 71.8 78.2 83.0
LLaVA-OneVision 72B 85.6 83.7 91.3 74.9 67.5 56.8 52.3 31.0 71.9 75.4 80.5 85.2
NVLM-D-1.0 78B 94.2 86.0 92.6 65.2 59.7 54.6 69.7 82.1 85.4
Llama 3.2 90B 92.3 85.5 90.1 57.3 60.3 39.5
3.2.1 Image Benchmarks
As presented in Table 8, we conduct comprehensive evaluations across a diverse range of image benchmarks: AI2D [47], ChartQA [48], DocVQA [49], InfographicVQA [50], MathVista [51], MMMU [52] (with zero-shot CoT), RealworldQA [53], SEED-Bench [54], TextVQA [55], and VQAv2 [56].
Our NVILA performs comparably to top open-source models in each size category, including Qwen2-VL [5], InternVL [3], and Pixtral. For general visual question answering tasks (ChartQA, DocVQA, InfoVQA, TextVQA, VQAv2, Seed), NVILA-8B and NVILA-15B achieve competitive or even better results compared to proprietary models (GPT-4o, Gemini). In science-related benchmarks (AI2D), NVILA-8B achieves state-of-the-art performance among open-source models. When scaling to 15B, NVILA demonstrates competitive performance with proprietary models.
Furthermore, on reasoning and knowledge benchmarks such as MMMU, RealworldQA, and MathVista, scores improve more when the model size increases. For benchmarks that require OCR capability such as TextVQA, AI2D, ChartQA, DocVQA, InfoVQA, 8B model can also do a great job. We also show a few qualitative examples in Figure. LABEL:fig:qualitative_example to demonstrate the OCR, reasoning, and multi-image capability of the NVILA model.
3.2.2 Video Benchmarks
We evaluate our models on a range of video understanding benchmarks [57, 58, 59, 60], spanning short videos of a few seconds to longer videos up to an hour in duration. Table 9 presents the performance of NVILA compared to baseline models [61, 62, 5, 4, 63, 19]. NVILA features long-context capability and can process up to 256 frames. With the scale-then-compress design, NVILA-8B achieves impressive results, setting new state-of-the-art performance across all benchmarks. Notably NVILA reaches performance levels comparable to GPT-4o mini with only 8B parameters and outperforms many larger models.
Table 9: Video benchmarks.
ActivityNet-QA LongVideoBench MLVU MVBench NExT-QA Video-MME
#F acc. score val test m-avg test mc w/o sub. w/ sub.
GPT-4o mini 56.5 58.8 64.8 68.9
GPT-4o 61.9 66.7 66.7 64.6 71.9 77.2
VILA-1.5 8B
LLaVA-NeXT-Video 7B 32 53.5 3.2 43.5 43.5 33.7 46.5
Video-XL 7B 2048 49.5 51.3 64.9 55.3 77.2 55.5 61.0
InternVL2 8B 64 54.6 64.0 65.8 56.3 59.3
LLaVA-OneVision 8B 32 56.6 56.5 64.7 56.7 79.4 58.2 61.5
Oryx-1.5 8B 128 56.3 67.5 67.6 81.8 58.8 64.2
LongVILA 7B 256 59.5 57.1 67.1 80.7 60.1 65.1
LongVU 7B 1fps 65.4 66.9 60.6
Qwen2-VL 8B 2fps 55.6 56.8 65.5 67.0 63.3 69.0
NVILA 8B 256 60.9 3.7 57.7 58.7 70.1 68.1 82.2 64.2 70.0
3.3 Efficiency Results
Figure 5: NVILA demonstrates superior inference efficiency over the Qwen2-VL model [5] for both image and video understanding tasks. We benchmark NVILA-7B against Qwen2-VL-7B. Qwen2-VL-7B is served by vLLM [64] for W4A16 LLM quantization, while NVILA is quantized and deployed with our specialized inference engine. Specifically, we ablate the efficiency gains achieved with different optimization techniques we introduced in NVILA. NVILA demonstrates 1.6-2.2$\times$ faster prefilling and up to 2.8$\times$ higher decoding throughput compared to Qwen2-VL.
NVILA achieves competitive performance on image and video benchmarks while maintaining efficiency through “scale-then-compress”. Architecturally, We initially scale up to native resolution (1–12× more tiles), then compress tokens by 2.4×, achieving higher accuracy with slightly more tokens than previous solutions. Dataset-wise, we curate a diverse 10M sample dataset, compress it using DeltaLoss, and prune to a high-quality 5M subset, consistently outperforms LLaVA-OneVision, which trained on 8M+ data. Besides, we integrate FP8 for acceleration, optimize learning rates for fine-tuning, and use W8A8 format to improve latency and throughput. These full-stack optimizations enable NVILA to train with fewer resources while achieving better performance, less memory usage, and faster inference.
We compare NVILA’s inference performance against Qwen2-VL [5] as shown in Figure 5. For a fair comparison, both models process video inputs by sampling 64 frames, with all experiments conducted on a single NVIDIA RTX 4090 GPU. Qwen2-VL is quantized to W4A16 and deployed with vLLM [64], a LLM/VLM serving engine with state-of-the-art inference speed. For NVILA, we quantize the LLM backbone to W4A16 and vision tower to W8A8. With our specialized inference engine, NVILA achieves up to 2.2$\times$ speedup in pre-filling stage and up to 2.8$\times$ higher decoding throughput over Qwen2-VL.
4 More Capabilities
4.1 Temporal Localization
Following LITA, we also add support for temporal localization in NVILA. We add discrete time tokens to indicate the timestamps in the video, and use the smoothed cross entropy loss to train the model. From the results in Table 10, we can clearly see that NVILA substantially outperforms all baselines for all metrics.
Table 10: Temporal localization. LITA results are from their original paper, while VILA-1.5 results are based on our reproduction. Our NVILA uses the same data mixture as VILA-1.5; the only difference is the backbone VLM.
ActivityNet-RTL
#Frames Mean IoU Precision@0.5
LITA 7B 100 24.1 21.1
LITA 13B 100 28.6 25.9
VILA-1.5 8B 256 32.1 29.3
NVILA 8B 256 34.8 32.1
Figure 6: Robotic navigation. NVILA deployed as a Vision-Language Navigation agent, navigating environments using language instructions and visual observations (Top: simulation, Bottom: real-world). The real-world setup features a Unitree Go2 robot equipped with a LiDAR sensor at the base of its head and an Intel RealSense Camera mounted on top. On the server side, an RTX 4090 GPU powers the NVILA-8B model, configured with an 8-frame context length for action generation.
4.2 Robotic Navigation
[…]
Table 11: Robotic navigation. All numbers are from NaVILA, except for those of NVILA. All models are provided with only RGB inputs. We refer the readers to NaVILA [8] for more details.
R2R Val-Unseen
Obs. NE $\downarrow$ OS $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq RGB 10.10 8.0 0.0 0.0
CMA RGB 9.55 10.0 5.0 4.0
NaVid 7B RGB 5.47 49.0 37.0 35.0
NVILA 8B RGB 5.43 60.4 53.3 48.8
4.3 Medical Application
NVILA also offers transformative potential in the medical domain. Such integration promises advancements in diagnostic accuracy, clinical decision-making, and data interpretation. The NVILA-M3 framework [11] introduces a novel approach by integrating multiple domain-expert models tailored to specific medical tasks, such as image segmentation and classification. These expert models are designed to extract and interpret intricate features that are otherwise difficult for general VLM’s to discern. By coupling these specialized models with a vision-language learning paradigm, NVILA-M3 achieves enhanced performance, facilitating the learning of nuanced relationships between visual inputs and their textual annotations. This integration not only improves task-specific outcomes but also sets a foundation for the development of more robust and context-aware VLMs in the healthcare domain. NVILA-M3 indicated that an overall improvement of 9% can be achieved via usage of expert models over existing SOTA, a few key results can be observed in Table. 12. This underscores the importance of leveraging domain expertise to bridge the gap between generalized AI capabilities and the demands of specialized applications, demonstrating the potential for VLMs to revolutionize fields where precision and specificity are paramount.
Table 12: Medical application. Performance of best M3 model on key benchmarks is shown. Task-specific SOTA baselines and datasets are described in the experiments section [11]. Metrics for VQA is accuracy, for report generation BLEU-4 & ROUGE and for classification F1 score have been utilized
VQA Report Gen. Classif.
Rad Path CXR CheXpert
Med-Gemini 78.8 83.3 20.5 28.3 48.3
VILA-M3 8B 84.7 91.0 21.1 32.0 61.6
NVILA 8B 85.5 92.9 22.8 32.8 61.1
Task-spfc. SOTA 84.2 91.7 15.4 30.6 51.5
5 Related Work
5.1 Visual Language Models
VLMs, especially proprietary ones, have advanced rapidly over the past two years. For example, OpenAI has upgraded from GPT-4V [67] to GPT-4o [12], achieving a 5–10% performance gain across image and video QA benchmarks. Google has extended the context length to 1M in Gemini Pro 1.5 [68], a significant improvement over Gemini 1.0 [69]. It now ranks at the top of the Video-MME leaderboard [60] for long video understanding. Anthropic has released Claude 3.5 [13], which demonstrates better benchmark scores than GPT-4o, showcasing notable improvements over Claude 3 [70]. Other proprietary models have similar advancements, such as Apple’s upgrade from MM1 to MM1.5 [71] and xAI’s upgrade from Grok-1.5 [53] to Grok-2 [72].
Meanwhile, open-source VLMs continue to evolve, improving at both the system/framework level [73] and the algorithm/recipe level [2], progressively narrowing the performance gap between proprietary and open-source models [19, 74, 75, 76, 5]. These recent advancements have led many open VLM models to claim performance levels comparable to, or even exceeding, leading proprietary models such as GPT-4V and GPT-4o. Some representative examples include InternVL2 [3], Qwen2-VL [5], LLaVA-OneVision [4], Llama 3.2 Vision [77], Molmo [30], NVLM [76], and MiniCPM-V [18].
Despite significant advancements in model performance, much less focus has been placed on enhancing the efficiency of training, inference, and fine-tuning for these models. This paper aims to explore how to develop VLMs that are not only highly accurate but also optimized for end-to-end efficiency.
5.2 Efficiency
Prior works such as [78, 79, 80, 62, 81, 82, 83, 84] have explored token reduction techniques in both spatial and temporal dimensions. However, none have focused on reducing the number of tokens for a frontier Vision-Language Model (VLM). For dataset pruning, promising approaches have been proposed for selecting pretraining data for Large Language Models (LLMs), such as domain-mixing [85], sample-wise data selection [27, 86], and theory-driven optimal selection [28]. In this work, we specifically focus on pruning supervised fine-tuning (SFT) datasets for VLMs. Regarding low-precision training, FP8 training [87, 88] has gained popularity for LLMs, yet no prior work has demonstrated its feasibility for VLMs without sacrificing accuracy. Techniques such as pruning, distillation, and quantization are commonly applied to LLMs. [89, 90] apply pruning/distillation to LLM. However, their application to VLMs presents an open question: Should an LLM be pruned or distilled first before integrating a vision encoder, or should the VLM itself be pruned or distilled after training? Similarly, quantization techniques like AWQ [41] and GPTQ [91] are well-documented for LLMs, and VILA [2] has shown that AWQ can be directly applied to VLMs. However, little attention has been given to quantizing vision encoders, which becomes critical when handling higher-resolution images or videos due to the increased computational demands. Parameter-efficient fine-tuning methods such as LoRA [92], DoRA [93], QLoRA [94], and GaLoRA [95] are widely used for LLMs to reduce memory requirements. However, for VLMs, which combine a vision encoder with an LLM, efficient fine-tuning techniques are still underexplored. Addressing this gap is crucial for advancing VLM fine-tuning with limited computational resources.
6 Conclusion
This paper introduces NVILA, a family of open VLMs designed to strike an optimal balance between efficiency and accuracy. By adopting the “scale-then-compress” paradigm, NVILA can efficiently process high-resolution images and long videos while maintaining high accuracy. We also systematically optimize its efficiency across the entire lifecycle, from training to fine-tuning to inference. NVILA delivers performance that matches or exceeds current leading VLMs, while being significantly more resource-efficient. Moreover, NVILA opens up new possibilities for applications such as temporal localization, robotic navigation, and medical imaging. We will make our models available soon. We hope NVILA can empower researchers and developers to fully unlock its potential across a wide range of applications and research domains.